OpenStudy (anonymous):

A statement Sn about the positive integers is given. Write statements Sk and Sk+1, simplifying Sk+1 completely. Sn: 1 ∙ 2 + 2 ∙ 3 + 3 ∙ 4 + . . . + n(n + 1) = [n(n + 1)(n + 2)]/3

OpenStudy (anonymous):

@satellite73 @sami-21 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

To find Sk, replace every 'n' with 'k'

jimthompson5910 (jim_thompson5910):

Do the same to find Sk+1

jimthompson5910 (jim_thompson5910):

Let me know what you get

OpenStudy (anonymous):

sk [k(k + 1)(k + 2)]/3 sk+1 [k+1(k+1 + 1)(k+1 + 2)]/3

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

More like this Sk: 1 ∙ 2 + 2 ∙ 3 + 3 ∙ 4 + . . . + k(k + 1) = [k(k + 1)(k + 2)]/3

jimthompson5910 (jim_thompson5910):

Sk+1: 1 ∙ 2 + 2 ∙ 3 + 3 ∙ 4 + . . . + (k+1)(k+1 + 1) = [(k+1)(k+1 + 1)(k+1 + 2)]/3 You'll need to simplify the above

OpenStudy (anonymous):

k(k + 1) = [k(k + 1)(k + 2)]/3